A stone of mass 2Kg is dropped freely from the top of a building . The stone covers 1/5th of the height of the building in the 5th second of its journey . The height of the building is

Grade:12

1 Answers

abhishek
9 years ago
let S1 distance in 4 sec
S1 = g(42)/2
let S2 = g(52)/2 distance covered in 5 sec
Now,
distance covered in 5th sec is the distance covered after completing 4sec till 5sec
so let hieght of building be h
then h/5 = S2 – S1
h = 45g
if g is taken as10 then h = 450m
if g = 9.8 then h = 441m